# SPECIAL RELATIVITY

### v11

@@ 10.1 The Potential of Earth-Based Electron Microscopes
We propose an experiment using commercial 200-kV scanning transmission electron microscope (STEM) as a particle source. One could use a vertically oriented optical setup. We suppose that the electron beam is imaged by a lens to a spot that lies in the image plane. Measurement of the spot shift with respect to the optical axis (which is parallel to the zenith-nadir axis) is equivalent to the determination of the angular displacement. Due to the high stability, the aberration increment could be observed  [^32] The dedicated STEM with high stability was made [47]. It was developed based on a 200-kV commercial instrument. The electron source was installed in the anti-seismic room. Temperature fluctuations in the room were within 0.2 ${}^{o}C$. The magnetic field in the room was less than $0.2mG$ near the source column. To prevent thermal drift, a shield box was placed over the high-voltage generator. The column was covered with rubber sheets to reduce the temperature fluctuations. The image drift measured during a period of 61 minutes was $\simeq 6~{}\mathrm{nm}$. The drift speed and direction during this period were approximately constant. The average drift speed was 0.12 nm/min. It is substantially smaller than that conventional commercial microscope, in which the drift speed is about 2 nm/min..
@@
A specific aspect of our case study needs further investigation. The proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. Obviously, it is important that observation should be recorded in the shortest possible time. In principle, records could be taken over a period of one day. The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth’s rotation axis 333333 The constant of diurnal aberration defines the scale of the aberration increment and is equal to $\theta_{d}=(1-1/1/\gamma)v_{e}/V\simeq 0.6$ $\mu\mathrm{rad}$. Here $\gamma\simeq 1.39$, $V/c=0.7$, $v_{e}/c\simeq 1,55$ $\mu\mathrm{rad}$ where $v_{e}=0.45$ km/s is the linear velocity of the earth rotation at equator.. In practice, the measurement procedure is complicated by a number of factors, Fig. 40. A correction has to be made for the observer’s position on the earth’s surface and the obliquity of the ecliptic $\delta=23^{o}$. The image appears to move in an ellipse. The shape of the aberration ellipse depends on the observer’s latitude $\phi$. For an observer on the equator, the ellipse degenerates into a line segment, and for an observer at the pole of the earth, the ellipse is a circle. At the value $l=2$ mm of the focal length  [^34] Electron-optical system of the STEM contains an electron gun and several magnetic lenses, stacked vertically to form a lens column. The illumination system of the instrument comprises the electron gun, together with two condenser lenses that focus the electrons onto the spacemen. Electrons entering the lens column appear to come from a virtual source. The first condenser lens (C1) is a strong magnetic lens, with a focal length $f_{1}$ of typically 2 mm. Using a virtual electron source as its object, C1 produces a real image. Because the lens is located $D_{1}=15$ cm below the object (virtual source), its object distance is 15 cm $\gg f_{1}$ and the image distance $\simeq f_{1}$. The second condenser lens (C2) is a weak magnetic lens ($f_{2}\simeq 10$ cm) that provides little magnification but allows the diameter of illumination at the specimen to vary continuously over a wide range. The distance between lens centers is typically 10 cm. Spacemen is located $D_{3}=25$ cm below the C2., the major axis of the aberration ellipse of the order of 80 ($\cos\delta)$ nm. The value of the minor axis depends on the observer’s position and is given by 80 ($\cos\delta\sin\phi))$ nm. When the latitude is $\phi=35^{o}$, the value of the scale of image shift per hour depends on the local time and varies from a value of about 10 nm/60 min to about 20 nm/60 min through the day  [^35] Experimental results (see Fig. 44 p.167 in [47]) shows that the image shift is quite close to the theoretical prediction $\simeq 10~{}\mathrm{nm}$. It is important to note that we used the typical focal length (2 mm) in our estimations. The details of the optical setup may influence the image shift. .
@@ 10.2 A Thought Experiment with a Point-Like Source
An incandescent lamp is a completely incoherent source. The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34]. In our case of interest, imaging system is situated in the far zone $D^{2}/(\lambda z)<1$. It is the source linear dimension $d$ (source diameter) that determines the coherent area of the observed radiation $(z\lambda/d)^{2}$. The condition for neglecting the transverse size of the source, $d$, formulated as the requirement that the change of the correlation function along the aperture be less than unity, that is $d<D$ 363636 In all cases of practical interest, telescopes are situated in the far zone of the distant stars. If the far zone approximation $z>d$ is satisfied, then the Van Cittert-Zernike theorem takes its simplest form. For example, the coherence area of the light emitted by the circular incoherent source is $A_{c}=4\lambda^{2}z^{2}/(\pi d^{2})$. The minimal coherent area of light observed on the earth’s surface from the nearest star has a diameter of about $\lambda z/d\simeq$ 10 m. This diameter is larger than a typical telescope aperture $D$. In this situation, a star can be considered as a point source and a star image is a point spread function in the image plane of a telescope. Our scaling model is based on the far zone approximation. We have only to change the distance and the source diameter. Assuming small parameter conditions $Dd/(z\lambda)<1$, $D^{2}/(z\lambda)<1$, the scaling approach yields the following requirements: $D^{2}/(z\lambda)<1$, $d<D$,.
Let us estimate the parameters for an Earth-based setup using an incandescent lamp. We aim to describe a series of thought experiments—hypothetical, yet physically plausible. At optical frequencies, the conditions required for the Fraunhofer approximation can be quite stringent. For instance, with a wavelength of $\lambda=0.3$ $\mu$m and an aperture width of $D=3$ cm, the observation distance must satisfy $z>3000$ m. Obviously, the required conditions are met in a number of important problems. No one has ever done such an experiment, but we know what would happen from the astronomical observations.
@@ 10.3 Practical Setup Using an Incoherent Optical Source
To perform an aberration of light measurement, the source linear dimension $d$ should be radically decreased. We assume that the order of magnitude of the dimensions of the ”point” source is about $\lambda$, where $\lambda$ is the optical radiation wavelength. Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms. In optics, the emission from such a ”point” source is described classically using Maxwell’s equations, while the emitting medium (considered as an ensemble of atoms) is treated quantum mechanically. Thanks to modern lithography techniques, it is relatively straightforward to create an unresolved point source at optical wavelengths, allowing for the generation of sufficiently bright radiation  [^37] It may be possible to use the masked spatially incoherent UV light source. One could use the light source onto a (micron diameter) hole of a specially designed mask. The mask should be placed at the distance $l<<d$, where $d$ is the transverse size of the spatially incoherent UV light source..
@@
Let us estimate the parameters for the optical setup shown in Fig.42. Suppose $a=1.5$ cm and $\lambda=0.3$ $\mu$m. We will use $NA=0.1$. The resolution analysis is reduced to the theoretical framework of standard imaging theory. One can take advantage of well-known resolution criteria like Rayleigh criteria i.e. $20.6\lambda/NA=$ 2 $\mu$m. One could use a vertically oriented optical setup. The constant of aberration defines the scale of the aberration increment, and is equal to $v/c=100$ $\mu$rad. Measurement of the spot shift with respect to the optical axis (which is parallel to the nadir direction) is equivalent to the determination of the angular displacement. As already pointed out, the proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. In principle, records could be taken over a period of one day. The image appears to move in an ellipse. The major axis of the ellipse is about 14 $\mu$m. The value of the minor axis depends on the observer’s position and is given by $14(\sin\phi)$ $\mu$m  [^38] The small image size—approximately 2 $\mu$m in the numerical case considered—places significant limitations on the use of a CCD as a detector. To address this, one promising alternative is photon detection using a microchannel plate photomultiplier (MCP-PMT). A feasible approach involves using a commercially available precision 2D translation stage with a resolution of 0.5 $\mu$m to control the position of a submicron aperture in an opaque screen. For instance, the aperture diameter could be set to 1 $\mu$m or smaller. The screen is mounted horizontally on the translation stage, which is positioned in the image plane. The MCP-PMT can then detect photons transmitted through the aperture. This setup offers the advantage of compactness, with the overall height of the optical configuration being approximately one meter..
@@
It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like oscillatory image shift within a period of one day. This argument gives the reason why our technique can be extended to allow for the reconstruction of the angle of aberration from a measurement of image shift of a finite size, incoherent source. The edge object has become useful in aberration shift detection. An alternative application of a ”point” source is an incoherent source with (physically) sharp edges. The edge width of the source is strictly related to the practical resolution achievable. For our particular example, the edge width should be about $1\mu$m.
@@ 11.1 Theory of a Laser with an Optical Resonator
The electric field $\vec{E}$ of an electromagnetic wave in a passive resonator satisfies the equation $\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. For simplicity, we shall assume that the field in the resonator is circularly polarized. Such an assumption does not violate the generality of the analysis because the polarization degeneracy takes place in the system under study. The electric field vector in the plane resonator is presented as an oscillation superposition with different longitudinal wave numbers, $E_{x}+iE_{y}=\sum_{m}\widetilde{E}_{m}(x,y,t)\exp(-i\omega_{m}t)\sin(K_{m}z+\delta)$, where $K_{m}=m\pi/l-i/(nl)$, $\omega_{m}=m\pi c/l$, $\delta=i/n$, $n$ is the (complex) refractive index of the mirror matter ($|n|\gg 1)$, $l$ is the distance between the mirrors, and $m$ is an integer number ($m\gg 1$). It is evident that this expression satisfies Leontovich’s boundary conduction  [^39] An example of introducing approximate boundary conditions is presented here. These conditions, originally derived by Leontovich, apply to surfaces of materials with a high refractive index modulus. A more comprehensive theoretical treatment can be found in the monograph [48]. on the mirror surface when $z=0$ and $z=l$. We assume the field change per one resonator pass to be small i.e. $|\partial\widetilde{E}_{m}/\partial t|l/c\ll|\widetilde{E}_{m}|$, which means that the laser frequency is close to the natural frequencies $\omega_{m}$ of the longitudinal resonator modes. Let us find a solution for the field amplitude $\widetilde{E}_{m}$ in the form $\widetilde{E}_{m}=\Phi(x,y)\exp(\Lambda t)$. Substituting the ansatz in the wave equation we get $\nabla_{\perp}\Phi+2i\omega_{m}[\Lambda/c^{2}+2/(nl)]\Phi=0$.
@@
In a plane resonator, the diffraction effects at the edges of the mirrors can be taken into account if one considers the space between the mirrors as a waveguide, and uses diffraction theory at the open end of the waveguide. If almost an integer number of half waves is laid along the length between the mirrors then at the open resonator end we may put boundary conditions on the amplitude function $\Phi(x,y)$, equivalent by its effect to the open waveguide end action  [^40] The analytical method is based on the introduction of specific, complex boundary conditions known as impedance boundary conditions of the resonance type [49]. : $\Phi+(1+i)\beta_{0}\sqrt{cl/(4\omega_{m})}\partial\Phi/\partial\eta=0$, where $\eta$ is the direction of the normal to the imaginary side surface of the resonator, and the parameter $\beta_{0}=0.824$  [^41] Interestingly, $\beta_{0}$ is related to one of the most renowned mathematical functions—the Riemann zeta function $\zeta$. Specifically, it is given by $-\zeta(1/2)/\sqrt{\pi}$; see [49].. The boundary conditions do not depend on the polarization which means that polarization degeneracy occurs in the system under study. Therefore, the problem concerning the open plane resonator excitation using equivalent boundary conditions can be reduced to the more usual classical formulation of the problem of closed resonator excitation.
[…]
@@ 11.2 Absent of Light Aberration from an Earth-Based Laser
The distinction between a ”plane wave” emitter and a laser source leads to a striking prediction in the theory of light aberration. Specifically, if a ”plane wave” emitter initially at rest with respect to the fixed stars begins to move, the apparent angular position of the emitter—as observed from the accelerated frame—undergoes a sudden shift by an angle of $-v/c$ (see Fig. 44). Understanding this phenomenon requires the framework of special relativity; classical theory is insufficient.
@@ 11.3 A Moving Laser Source in an Inertial Frame of Reference
When discussing light aberration, it is important to distinguish between the effect from a ”plane wave” emitter and that from the laser source. Consider a ”plane wave” emitter that is accelerated from rest to velocity $v$ in a direction perpendicular to its optical axis  [^42] It is assumed that after acceleration, the laser moves at constant velocity and operates in a steady-state regime.. Suppose an observer, at rest with respect to the inertial frame of reference, measures the direction of light transport. How does the light beam from the moving ”plane wave” emitter appear to this observer? The inertial observer would find an angular displacement of $\theta_{a}=v/c$, as illustrated in Fig. 46.

### v12

@@ 10.1 The Potential of Earth-Based Electron Microscopes
We propose an experiment using commercial 200-kV scanning transmission electron microscope (STEM) as a particle source. One could use a vertically oriented optical setup. We suppose that the electron beam is imaged by a lens to a spot that lies in the image plane. Measurement of the spot shift with respect to the optical axis (which is parallel to the zenith-nadir axis) is equivalent to the determination of the angular displacement. Due to the high stability, the aberration increment could be observed.  [^34] The dedicated STEM with high stability was made [47] It was developed based on a 200-kV commercial instrument. The electron source was installed in the anti-seismic room. Temperature fluctuations in the room were within 0.2 $^{o}C$. The magnetic field in the room was less than $0.2mG$ near the source column. To prevent thermal drift, a shield box was placed over the high-voltage generator. The column was covered with rubber sheets to reduce the temperature fluctuations. The image drift measured during a period of 61 minutes was $\simeq 6~\mathrm{nm}$. The drift speed and direction during this period were approximately constant. The average drift speed was 0.12 nm/min. It is substantially smaller than that conventional commercial microscope, in which the drift speed is about 2 nm/min.
@@
A specific aspect of our case study needs further investigation. The proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. Obviously, it is important that observation should be recorded in the shortest possible time. In principle, records could be taken over a period of one day. The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth’s rotation axis 353535 The constant of diurnal aberration defines the scale of the aberration increment and is equal to $\theta_{d}=(1-1/1/\gamma)v_{e}/V\simeq 0.6$ $\mu\mathrm{rad}$. Here $\gamma\simeq 1.39$, $V/c=0.7$, $v_{e}/c\simeq 1,55$ $\mu\mathrm{rad}$ where $v_{e}=0.45$ km/s is the linear velocity of the earth rotation at equator.. In practice, the measurement procedure is complicated by a number of factors, Fig. 40. A correction has to be made for the observer’s position on the earth’s surface and the obliquity of the ecliptic $\delta=23^{o}$. The image appears to move in an ellipse. The shape of the aberration ellipse depends on the observer’s latitude $\phi$. For an observer on the equator, the ellipse degenerates into a line segment, and for an observer at the pole of the earth, the ellipse is a circle. At the value $l=2$ mm of the focal length, the major axis of the aberration ellipse of the order of 80 ($\cos\delta)$ nm.  [^36] Electron-optical system of the STEM contains an electron gun and several magnetic lenses, stacked vertically to form a lens column. The illumination system of the instrument comprises the electron gun, together with two condenser lenses that focus the electrons onto the spacemen. Electrons entering the lens column appear to come from a virtual source. The first condenser lens (C1) is a strong magnetic lens, with a focal length $f_{1}$ of typically 2 mm. Using a virtual electron source as its object, C1 produces a real image. Because the lens is located $D_{1}=15$ cm below the object (virtual source), its object distance is 15 cm $\gg f_{1}$ and the image distance $\simeq f_{1}$. The second condenser lens (C2) is a weak magnetic lens ($f_{2}\simeq 10$ cm) that provides little magnification but allows the diameter of illumination at the specimen to vary continuously over a wide range. The distance between lens centers is typically 10 cm. Spacemen is located $D_{3}=25$ cm below the C2. The value of the minor axis depends on the observer’s position and is given by 80 ($\cos\delta\sin\phi))$ nm. When the latitude is $\phi=35^{o}$, the value of the scale of image shift per hour depends on the local time and varies from a value of about 10 nm/60 min to about 20 nm/60 min through the day.  [^37] Experimental results (see Fig. 44 p.167 in [47]) shows that the image shift is quite close to the theoretical prediction $\simeq 10~\mathrm{nm}$. It is important to note that we used the typical focal length (2 mm) in our estimations. The details of the optical setup may influence the image shift.
@@ 10.2 A Thought Experiment with a Point-Like Source
An incandescent lamp is a completely incoherent source. The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34]. In our case of interest, imaging system is situated in the far zone $D^{2}/(\lambda z)<1$. It is the source linear dimension $d$ (source diameter) that determines the coherent area of the observed radiation $(z\lambda/d)^{2}$. The condition for neglecting the transverse size of the source, $d$, formulated as the requirement that the change of the correlation function along the aperture be less than unity, that is $d<D$. 383838 In all cases of practical interest, telescopes are situated in the far zone of the distant stars. If the far zone approximation $z>d$ is satisfied, then the Van Cittert-Zernike theorem takes its simplest form. For example, the coherence area of the light emitted by the circular incoherent source is $A_{c}=4\lambda^{2}z^{2}/(\pi d^{2})$. The minimal coherent area of light observed on the earth’s surface from the nearest star has a diameter of about $\lambda z/d\simeq$ 10 m. This diameter is larger than a typical telescope aperture $D$. In this situation, a star can be considered as a point source and a star image is a point spread function in the image plane of a telescope. Our scaling model is based on the far zone approximation. We have only to change the distance and the source diameter. Assuming small parameter conditions $Dd/(z\lambda)<1$, $D^{2}/(z\lambda)<1$, the scaling approach yields the following requirements: $D^{2}/(z\lambda)<1$, $d<D$.
[…]
@@ 10.3 Practical Setup Using an Incoherent Optical Source
To perform an aberration of light measurement, the source linear dimension $d$ should be radically decreased. We assume that the order of magnitude of the dimensions of the ”point” source is about $\lambda$, where $\lambda$ is the optical radiation wavelength. Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms. In optics, the emission from such a ”point” source is described classically using Maxwell’s equations, while the emitting medium (considered as an ensemble of atoms) is treated quantum mechanically. Thanks to modern lithography techniques, it is relatively straightforward to create an unresolved point source at optical wavelengths, allowing for the generation of sufficiently bright radiation.  [^39] It may be possible to use the masked spatially incoherent UV light source. One could use the light source onto a (micron diameter) hole of a specially designed mask. The mask should be placed at the distance $l<<d$, where $d$ is the transverse size of the spatially incoherent UV light source.
@@
[…]
@@
[…]
@@ 11.1 Theory of a Laser with an Optical Resonator
The electric field $\vec{E}$ of an electromagnetic wave in a passive resonator satisfies the equation $\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. For simplicity, we shall assume that the field in the resonator is circularly polarized. Such an assumption does not violate the generality of the analysis because the polarization degeneracy takes place in the system under study. The electric field vector in the plane resonator is presented as an oscillation superposition with different longitudinal wave numbers, $E_{x}+iE_{y}=\sum_{m}\widetilde{E}_{m}(x,y,t)\exp(-i\omega_{m}t)\sin(K_{m}z+\delta)$, where $K_{m}=m\pi/l-i/(nl)$, $\omega_{m}=m\pi c/l$, $\delta=i/n$, $n$ is the (complex) refractive index of the mirror matter ($|n|\gg 1)$, $l$ is the distance between the mirrors, and $m$ is an integer number ($m\gg 1$). It is evident that this expression satisfies Leontovich’s boundary conduction on the mirror surface when $z=0$ and $z=l$.  [^41] An example of introducing approximate boundary conditions is presented here. These conditions, originally derived by Leontovich, apply to surfaces of materials with a high refractive index modulus. A more comprehensive theoretical treatment can be found in the monograph [48]. We assume the field change per one resonator pass to be small i.e. $|\partial\widetilde{E}_{m}/\partial t|l/c\ll|\widetilde{E}_{m}|$, which means that the laser frequency is close to the natural frequencies $\omega_{m}$ of the longitudinal resonator modes. Let us find a solution for the field amplitude $\widetilde{E}_{m}$ in the form $\widetilde{E}_{m}=\Phi(x,y)\exp(\Lambda t)$. Substituting the ansatz in the wave equation we get $\nabla_{\perp}\Phi+2i\omega_{m}[\Lambda/c^{2}+2/(nl)]\Phi=0$.
@@
In a plane resonator, the diffraction effects at the edges of the mirrors can be taken into account if one considers the space between the mirrors as a waveguide, and uses diffraction theory at the open end of the waveguide. If almost an integer number of half waves is laid along the length between the mirrors then at the open resonator end we may put boundary conditions on the amplitude function $\Phi(x,y)$, equivalent by its effect to the open waveguide end action : $\Phi+(1+i)\beta_{0}\sqrt{cl/(4\omega_{m})}\partial\Phi/\partial\eta=0$, where $\eta$ is the direction of the normal to the imaginary side surface of the resonator, and the parameter $\beta_{0}=0.824$.  [^42] The analytical method is based on the introduction of specific, complex boundary conditions known as impedance boundary conditions of the resonance type [49]. Interestingly, $\beta_{0}$ is related to one of the most renowned mathematical functions—the Riemann zeta function $\zeta$. Specifically, it is given by $\zeta(1/2)/\sqrt{\pi}$; see [49]. The boundary conditions do not depend on the polarization which means that polarization degeneracy occurs in the system under study. Therefore, the problem concerning the open plane resonator excitation using equivalent boundary conditions can be reduced to the more usual classical formulation of the problem of closed resonator excitation.
Let us consider a resonator equipped with plane circular mirrors with radius $R$. The distance between the mirrors is equal to $l$. The system is azimuthally symmetric relative to the resonator axis which coincides with the $z$ axis of the polar coordinates. We shall seek the solution of the wave equation in the following form: $\Phi(\vec{r})=\Phi_{\nu}(r)\cos(\nu\phi),\Phi_{\nu}(r)\sin(\nu\phi)$, where $\nu$ is an integer.The eigenfunctions of the resonator are the solution of the homogeneous equation $r^{2}d^{2}\Phi_{\nu j}/dr^{2}+rd\Phi_{\nu j}/dr+(k^{2}_{\nu j}r^{2}-\nu^{2})\Phi_{\nu j}=0$, satisfying the impedance boundary conditions. The eigenfunctions $\Phi_{\nu j}$ are orthogonal and form a complete system. In the first approximation for small parameter $M=1/\sqrt{N}$, where $N=\omega_{m}R^{2}/(cl)\gg 1$ can be referred to as the Fresnel number, the functions $\Phi_{\nu j}$ have the form $\Phi_{\nu}(r)=J_{\nu}(k_{\nu j}r)/J_{\nu+1}(\mu_{\nu j})$, where $\mu_{\nu j}$ is thejjth root of the Bessel function of order $\nu$, $k_{\nu j}=\mu_{\nu j}(1-\Delta)/R$, $\Delta=(1+i)\beta_{0}M/2$. Each mode for $\nu>0$ is four times degenerate. Firstly, for each value of the azimuthal number $\nu>0$ there are two linear independent functions $\propto\cos(\nu\phi)$ and $\propto\sin(\nu\phi)$. Secondly, because the boundary conditions do not depend on the polarization, two different field polarizations exist for each linear independent eigenfunction.
@@ 11.2 Absent of Light Aberration from an Earth-Based Laser
The distinction between a ”plane wave” emitter and a laser source leads to a striking prediction in the theory of light aberration. Specifically, if a ”plane wave” emitter initially at rest with respect to the fixed stars begins to move, the apparent angular position of the emitter—as observed from the accelerated frame—undergoes a sudden shift by an angle of v/c-v/c (see Fig. 44). Understanding this phenomenon requires the framework of special relativity; classical theory is insufficient.
@@ 11.3 A Moving Laser Source in an Inertial Frame of Reference
When discussing light aberration, it is important to distinguish between the effect from a ”plane wave” emitter and that from the laser source. Consider a ”plane wave” emitter that is accelerated from rest to velocity $v$ in a direction perpendicular to its optical axis.  [^43] It is assumed that after acceleration, the laser moves at constant velocity and operates in a steady-state regime. Suppose an observer, at rest with respect to the inertial frame of reference, measures the direction of light transport. How does the light beam from the moving ”plane wave” emitter appear to this observer? The inertial observer would find an angular displacement of $\theta_{a}=v/c$, as illustrated in Fig. 46.